OpenStudy (anonymous):

Determine without graphing whether the given quadratic function has a maximum value or a minimum value and then find the value f(x)=-3x^2+12x-3

OpenStudy (phi):

the trick to this one is the sign of the coefficient on the x^2 term if it is negative, then for "big x" you will get a negative number the graph looks like a "frown" if the number is positive, you get a "smile" hopefully you know which has a max and which has a min value

OpenStudy (anonymous):

yes, but whats next? okay I know its maximum, I guess I'm trying to figure out how they got 9

OpenStudy (phi):

the "vertex" of a parabola occurs at x= -b/2a where a,b,c are the numbers: ax^2 + bx + c in your case, a= -3 and b is +12 (c is -3 but we don't care for this problem) x= -b/(2a) means replace b with 12 and a with -3 to get x= - 12/(2*-3) = -12/-6 = +2 the y value is found by replacing x with 2 in the equation

OpenStudy (anonymous):

okay thanks, I understand now, once again Thanks!!!

OpenStudy (phi):

you can take a look at http://www.khanacademy.org/math/trigonometry/polynomial_and_rational/solving_graphing_quadratics/v/quadratic-functions-3 as a quick refresher. He explains this stuff pretty well.